Grade 12Physical Chemistry

The pKa of acetic acid and pKb of ammonium hydroxide are 4.76 and 4.75 respectively. Calculate the pH of ammonium acetate solution. ?

sudhanshu
12 Years agoGrade 12

2 Answers

mgayathri
10 Years ago
when a weak acid and a weak baseare mixed to form a solution,the ph of the solution is determined by following formula
ph=7+1/2[pka-pkb]
    =7+1/2[4.76-4.75]
     =7+1/2[0.01]
     =7+0.005
ph=7.005
Kushagra Madhukar
6 Years ago
Dear student,
Please find the attached solution to your problem.
 
We know, when a weak acid and a weak base are mixed to form a solution, the resulting ph of the solution is determined by following formula
pH = 7 + 0.5[pka-pkb]
    = 7 + 0.5[4.76 - 4.75]
    = 7 + 0.5[0.01]
    = 7 + 0.005
pH = 7.005
 
Hope it helps.
Thanks and regards,
Kushagra